A lady is 24 years older then her daughter. What will be the daughter's age when it is one-third of her mother's ?
Answer by nerdybill(7384)

A lady is 24 years older then her daughter. What will be the daughter's age when it is one-third of her mother's ?
.
Let d = daughter's age
then
d+24 = mother's age
.
d = (1/3)(d+24)
3d = d + 24
2d = 24
d = 12 years old